Suppose a biologist studying the mechanical limitations of growth among different species of tulips monitors a national preserve. He collects data on the heights of 10 different types of tulips in the reserve and rounds each height to the nearest centimeter. 25, 21,26, 24,29,33,29,25, 19,24 Compute the first quartile (Q1), the third quartile (Q3), and the interquartile range (IQR) of the data set. 022.5 cm 0329 cm IQR 6.5 cm
Answer #1

Q1 = 24

Q3 = 29

IQR = 5

Explanation:

First we have to arrange the numbers in assending order

19, 21, 24, 24, 25, 25, 26, 29, 29, 33

The lower half of the set of number is

19, 21, 24, 24, 25

Q1 is the median of the lower half of the data

So, Q1 = 24

The upper half of the data is

25, 26, 29, 29, 33

Q3 is the median of the upper half of the data

Q3 = 29

IQR = Q3 - Q1 = 29 - 24 = 5